# Volumetric Functional Maps

### v1

@@ 4.2. Spectral Connectivity Transfer
Knowing the ground truth correspondence $\pi:\partial\mathcal{M}\to\partial\mathcal{N}$, we can transfer the surface coordinates $x_{\partial\mathcal{N}},y_{\partial\mathcal{N}},z_{\partial\mathcal{N}}$ of $\partial\mathcal{N}$ onto $\partial\mathcal{M}$ as $T_{\pi}(x_{\partial\mathcal{N}}),T_{\pi}(y_{\partial\mathcal{N}}),T_{\pi}(z_{\partial\mathcal{N}})$. We then consider the LBO eigenbasis $\mathbf{\Phi}_{\mathcal{M}}$ and its restriction $\mathbf{\Phi}_{\partial\mathcal{M}}$ to the surface. Following a reasoning analogous to the discussion in the previous paragraph, we have that
(9) $\begin{gathered}(T_{\pi^{\prime}}(x_{\mathcal{N}}))_{\partial}\approx(\mathbf{\Phi}_{\mathcal{M}}\mathbf{\Phi}_{\mathcal{M}}^{\dagger}T_{\pi^{\prime}}(x_{\mathcal{N}}))_{\partial}=\mathbf{\Phi}_{\partial\mathcal{M}}\mathbf{\Phi}_{\mathcal{M}}^{\dagger}T_{\pi^{\prime}}(x_{\mathcal{N}})\,,\\ T_{\pi}(x_{\partial\mathcal{N}})\approx\mathbf{\Phi}_{\partial\mathcal{M}}\mathbf{\Phi}_{\partial\mathcal{M}}^{\dagger}T_{\pi}(x_{\partial\mathcal{N}})\,.\end{gathered}$
Since $(T_{\pi^{\prime}}(x_{\mathcal{N}}))_{\partial}=T_{\pi}(x_{\partial\mathcal{N}})$, then it must also be $\mathbf{\Phi}_{\mathcal{M}}^{\dagger}T_{\pi^{\prime}}(x_{\mathcal{N}})\approx\mathbf{\Phi}_{\partial\mathcal{M}}^{\dagger}T_{\pi}(x_{\partial\mathcal{N}})$. Finally, the coordinates on the interior of $\mathcal{N}$ can be extrapolated by reconstructing with the full basis:
(10) $\begin{gathered}T_{\pi^{\prime}}(x_{\mathcal{N}})=\mathbf{\Phi}_{\mathcal{M}}\mathbf{\Phi}_{\partial\mathcal{M}}^{\dagger}T_{\pi}(x_{\partial\mathcal{N}})\,,\\ T_{\pi^{\prime}}(y_{\mathcal{N}})=\mathbf{\Phi}_{\mathcal{M}}\mathbf{\Phi}_{\partial\mathcal{M}}^{\dagger}T_{\pi}(y_{\partial\mathcal{N}})\,,\\ T_{\pi^{\prime}}(z_{\mathcal{N}})=\mathbf{\Phi}_{\mathcal{M}}\mathbf{\Phi}_{\partial\mathcal{M}}^{\dagger}T_{\pi}(z_{\partial\mathcal{N}})\,.\end{gathered}$
Concrete examples of using these two approaches for reliable connectivity transfer are depicted in Figure 8. For this specific test, we extended the LBO eigenbasis with the Coordinates Manifold Harmonics (CMH) basis (Marin et al., 2019), which also includes three additional functions encoding the extrinsic per vertex $xyz$ coordinates and is known to be better suited to reconstruct extrinsic information. Additional experiments on the whole dataset, considering both the LBO and CMH eigenbasis are reported in Section 5.2.
@@ 4.3. Volume-Aware Surface Correspondence
Besides purely volumetric applications, we observe that a good functional map among volumes also induces a good functional map among their external surfaces. Intuitively, a volumetric shape is a richer structure than a surface, yielding more information both locally and globally. Therefore, we guess that functional maps among tetrahedral meshes are more accurate and informative than maps among triangular meshes.
@@
Once we have the tetrahedral meshes $\mathcal{M},\mathcal{N}$, we can compute the functional map $\mathbf{C}$ relating their eigenbases via $\mathbf{\Phi}_{\mathcal{M}}\mathbf{C}=T_{\pi}(\mathbf{\Phi}_{\mathcal{N}})$, and we can use that same functional map for defining the surface relation $(\mathbf{\Phi}_{\mathcal{M}})_{\partial}\mathbf{C}=(T_{\pi}(\mathbf{\Phi}_{\mathcal{N}}))_{\partial}$. Therefore, the surface correspondence can be extracted via the nearest neighbor search $\mathrm{NN}(\mathbf{\Phi}_{\partial\mathcal{M}}\mathbf{C},\mathbf{\Phi}_{\partial\mathcal{N}})$.
@@ 5. Results and Applications
We implemented our software pipeline in Matlab, starting from the reference implementations for the functional map estimation with descriptors preservation (Nogneng and Ovsjanikov, 2017) and ZoomOut (Melzi et al., 2019b), extending these tools to enable volume spectral processing, as detailed in Section 4 and in the remainder of this section.
@@ 5.1. Volumetric Functional Maps
Figure 6. Number of flipped tetrahedra obtained by transferring the mesh connectivity considering a growing number of basis functions. From left to right: coordinate transfer using the LBO basis; coordinate transfer using the CMH basis; coordinate extrapolation using the LBO basis; coordinate extrapolation using the CMH basis. The smaller colored lines show the trend for each pair of shapes in our reference dataset (Su et al., 2019). The bold black line is the average trend across the entire dataset.
@@
Table 1. Average percentage of flips for the methods from Figure 6 at varying of the basis size. The last column also report the average time (in seconds) required for the computation of the largest basis considered.
@@ 5.2. Connectivity Transfer
Representing volume maps using two tetrahedral meshes that share the same connectivity but have different vertex coordinates is a common choice for many practical applications involving digital fabrication (Liu et al., 2024; Zhang et al., 2022; Etienne et al., 2019), medicine (Abulnaga et al., 2021) and hexahedral meshing (Pietroni et al., 2022; Brückler and Campen, 2023). Our framework can be used to transfer the connectivity of a given tetrahedral mesh onto a target domain, either by using a volumetric functional map or by extrapolating the connectivity from its outer surface, as detailed in Section 4.2.
Figure 7. Left: one of the 1405 failure cases of the untangling algorithm proposed by Garanzha et al. (2021), which emerged from the challenging large scale validation described in (Nigolian et al., 2023, §6.3). Right: computing an initial connectivity transfer with our tool to warm start the same algorithm allows to retain a fully bijective map, thus improving the overall robustness.
@@
If the whole spectrum is considered, the LBO basis spans the entire functional space of the mesh and the functional map yields one-to-one correspondences (Rodola et al., 2017). Computing the full spectrum is computationally expensive, hence only a fraction of the eigenfunctions are typically computed. We test our prototypes on the dataset of 40 meshes released by Su  et al. (2019) and empirically verify that the amount of inverted elements introduced in the piece-wise map gracefully decays to zero for growing numbers of eigenfunctions (Figure 6). On average, already using $5\%$ of the spectrum the number of flipped tetrahedra is below $5\%$. It goes below $1\%$ if $20\%$ of the spectrum is considered (Table 1). We consider both the LBO spectrum and its extension equipped with three additional functions that encode the extrinsic embedding (CMH). As shown in (Marin et al., 2019; Melzi et al., 2020) CMH is better suited to transfer discrete signals such as the surface mesh connectivity. This advantage also manifests on volumes, as reflected in the second and fourth columns in Figure 6 and in the second and fourth lines of Table 1. For efficiency, the average running times for computing the larger bases are reported in the last column of the same table. As can be noticed the extrapolation algorithm is roughly 2×2$\times$2 × faster than volumetric transfer, because in this case only the volumetric spectrum of a single mesh must be computed.
Our results on connectivity transfer suggest that, even when the computational resources are limited, our tool could be used to reliably warm start mesh untangling methods such as (Garanzha et al., 2021; Du et al., 2020), which will then be required to sanitize the map by removing only a fraction of flipped tetrahedra, improving the overall robustness (Figure 7). Interestingly, if the mesh resolution is sufficiently low and the computation of a higher percentage of the spectrum is computationally feasible, we can directly devise a fully injective piecewise linear map, as shown in Figure 8, where $25\%$ of the spectrum is considered.
@@
5.3. Segmentation Transfer
The functional maps framework proved to be particularly successful in the area of statistical shape analysis (Stegmann and Gomez, 2002). In particular, the study of medical data primarily benefits from accurate solutions for analyzing large collections of shapes (Heimann and Meinzer, 2009), and the literature proves that the functional maps framework can be successfully applied for medical applications (Melzi et al., 2016; Magnet et al., 2023). A standard approach for analyzing collections of shapes is to determine or build a template surface which acts as a median shape for comparing the entire dataset (Huang et al., 2019). However, existing approach are limited to the analysis of surfaces, while often medical data are better represented as 3D volumes, such in the case of MRI scans (Rinck, 2019).
@@ 5.4. Shape Matching
We consider both the outer surfaces of our reference ground truth volumetric dataset (which we will refer to as VOL) (Su et al., 2019) and well established datasets for shape matching, such as SHREC’19 (Melzi et al., 2019a), SHREC’20 (Dyke et al., 2020), and TOPKIDS (Lähner et al., 2016), which we all tetrahedralize using fTetWild (Hu et al., 2020). Overall, these datasets stress shape matching algorithms in different ways, exhibiting geometric and topological noise (SHREC’19), strongly non-isometric matches (SHREC’20) and topological changes (SHREC’20 and TOPKIDS). For SHREC’20, we discard the two partial meshes shown in Figure 10 because the inside/outside system of fTetWild was not able to correctly reproduce a volumetric mesh that allows a meaningful comparison with surface methods. For the same reason, we are not able to consider alternative datasets such as SMAL (Zuffi et al., 2017), which contains numerous self-intersections that yield topological changes during tetrahedralization. Attempts to sanitize these models with various combinations of geometry processing tools (Portaneri et al., 2022; Hang, 2015; Livesu, 2019; Diazzi and Attene, 2021; Diazzi et al., 2023) ended with no luck. Indeed, our difficulties in creating a dataset that allows to meaningfully compare surface-based and volume-based shape matching methods raises an open research challenge that will hopefully be addressed by the community in the near future (Section 6).
@@ 6. Conclusions
First and foremost, a theoretical study of what properties are invariant under a volume functional map must be carried out. Despite our empirical studies show that volumetric maps are more accurate than their surface counterparts, theoretically justifying this behavior would provide useful insights for the development of new solutions.

### v2

@@ 4.2. Spectral Connectivity Transfer
Knowing the ground truth correspondence $\pi:\partial\mathcal{M}\to\partial\mathcal{N}$, we can transfer the surface coordinates $x_{\partial\mathcal{N}},y_{\partial\mathcal{N}},z_{\partial\mathcal{N}}$ of $\partial\mathcal{N}$ onto $\partial\mathcal{M}$ as $T_{\pi}(x_{\partial\mathcal{N}}),T_{\pi}(y_{\partial\mathcal{N}}),T_{\pi}(z_{\partial\mathcal{N}})$. We then consider the LBO eigenbasis $\Phi_{\mathcal{M}}$ and its restriction $\Phi_{\partial\mathcal{M}}$ to the surface. Following a reasoning analogous to the discussion in the previous paragraph, we have that
(9) $\begin{gathered}(T_{\pi^{\prime}}(x_{\mathcal{N}}))_{\partial}\approx(\Phi_{\mathcal{M}}\Phi_{\mathcal{M}}^{\dagger}T_{\pi^{\prime}}(x_{\mathcal{N}}))_{\partial}=\Phi_{\partial\mathcal{M}}\Phi_{\mathcal{M}}^{\dagger}T_{\pi^{\prime}}(x_{\mathcal{N}})\,,\\ T_{\pi}(x_{\partial\mathcal{N}})\approx\Phi_{\partial\mathcal{M}}\Phi_{\partial\mathcal{M}}^{\dagger}T_{\pi}(x_{\partial\mathcal{N}})\,.\end{gathered}$
Since $(T_{\pi^{\prime}}(x_{\mathcal{N}}))_{\partial}=T_{\pi}(x_{\partial\mathcal{N}})$, then it must also be $\Phi_{\mathcal{M}}^{\dagger}T_{\pi^{\prime}}(x_{\mathcal{N}})\approx\Phi_{\partial\mathcal{M}}^{\dagger}T_{\pi}(x_{\partial\mathcal{N}})$. Finally, the coordinates on the interior of $\mathcal{N}$ can be extrapolated by reconstructing with the full basis:
(10) $\begin{gathered}T_{\pi^{\prime}}(x_{\mathcal{N}})=\Phi_{\mathcal{M}}\Phi_{\partial\mathcal{M}}^{\dagger}T_{\pi}(x_{\partial\mathcal{N}})\,,\\ T_{\pi^{\prime}}(y_{\mathcal{N}})=\Phi_{\mathcal{M}}\Phi_{\partial\mathcal{M}}^{\dagger}T_{\pi}(y_{\partial\mathcal{N}})\,,\\ T_{\pi^{\prime}}(z_{\mathcal{N}})=\Phi_{\mathcal{M}}\Phi_{\partial\mathcal{M}}^{\dagger}T_{\pi}(z_{\partial\mathcal{N}})\,.\end{gathered}$
A visual representation of our pipeline is shown in Figure 4.
Concrete examples of reliable connectivity transfer using both a functional transfer of interior coordinates and their extrapolation from the spectral embedding of the surface coordinates are depicted in Figure 9 (resp. Transfer and Extrapolation). For this specific test, we extended the LBO eigenbasis with the Coordinates Manifold Harmonics (CMH) basis (Marin et al., 2019), which also includes three additional functions encoding the extrinsic per vertex $xyz$ coordinates and is known to be better suited to reconstruct extrinsic information. Additional experiments on the whole dataset, considering both the LBO and CMH eigenbasis are reported in Section 5.2.
@@ 4.3. Volume-Aware Surface Correspondence
Besides purely volumetric applications, we observe that a good functional map among volumes also induces a good functional map among their external surfaces. Intuitively, a volumetric shape is a richer structure than a surface, yielding more information both locally and globally. Therefore, we guess that functional maps among tetrahedral meshes are more accurate and informative than maps among triangular meshes. This idea is grounded from the results of Raviv et al. (2010), who have shown that using volume-informed shape descriptors has a positive effects in computing correspondences between surfaces.
@@
Once we have the tetrahedral meshes $\mathcal{M},\mathcal{N}$, we can compute the functional map $\mathbf{C}$ relating their eigenbases via $\Phi_{\mathcal{M}}\mathbf{C}=T_{\pi}(\Phi_{\mathcal{N}})$, and we can use that same functional map for defining the surface relation $(\Phi_{\mathcal{M}})_{\partial}\mathbf{C}=(T_{\pi}(\Phi_{\mathcal{N}}))_{\partial}$. Therefore, the surface correspondence can be extracted via the nearest neighbor search $\mathrm{NN}(\Phi_{\partial\mathcal{M}}\mathbf{C},\Phi_{\partial\mathcal{N}})$.
@@ 5. Results and Applications
We implemented our software pipeline in Matlab, starting from the reference implementations for the functional map estimation with descriptors preservation (Nogneng and Ovsjanikov, 2017) and ZoomOut (Melzi et al., 2019b), extending these tools to enable volume spectral processing, as detailed in Section 4 and in the remainder of this section. Our code is available at anonymous.
@@ 5.1. Volumetric Functional Maps
Figure 7. Number of flipped tetrahedra obtained by transferring the mesh connectivity considering a growing number of basis functions. From left to right: coordinate transfer using the LBO basis; coordinate transfer using the CMH basis; coordinate extrapolation using the LBO basis; coordinate extrapolation using the CMH basis. The smaller colored lines show the trend for each pair of shapes in our reference dataset (Su et al., 2019). The bold black line is the average trend across the entire dataset. The two pairs that take longer to converge with the LBO basis functions (yellow and orange lines) are the ones where source and target differ the most and are strongly non-isometric. This explains why using only the intrinsic information from the LBO eigenbasis makes the method converge slowly and adding extrinsic information from CMH aligns the pairs with the other results.
@@
Table 1. Average percentage of flips for the methods from Figure 7 at varying of the basis size. The last column also reports the average time (in seconds) required for the computation of the largest basis considered.
@@
Figure 8. Left: two of the 1405 failure cases of the untangling algorithm proposed by Garanzha et al. (2021), which emerged from the challenging large scale validation described in (Nigolian et al., 2023, §6.3). Right: computing an initial connectivity transfer with our tool to warm start the same algorithm allows to retain a fully bijective map, thus improving the overall robustness.
@@ 5.2. Connectivity Transfer
Representing volume maps using two tetrahedral meshes that share the same connectivity but have different vertex coordinates is a common choice for many practical applications involving digital fabrication (Liu et al., 2024; Zhang et al., 2022; Etienne et al., 2019), medicine (Abulnaga et al., 2021) and hexahedral meshing (Pietroni et al., 2022; Brückler and Campen, 2023). As detailed in Section 4.2, our framework can be used to transfer the connectivity of a given tetrahedral mesh onto a target domain, either by using a volumetric functional map (§4.2 Functional Maps for Connectivity Transfer) or by extrapolating the connectivity from its outer surface (§4.2 Spectral Representation of Volumes).
@@
If the whole spectrum is considered, the LBO basis spans the entire functional space of the mesh and the functional map yields one-to-one correspondences (Rodola et al., 2017). Computing the full spectrum is computationally expensive, hence only a fraction of the eigenfunctions are typically computed. We test our prototypes on the dataset of 40 meshes released by Su  et al. (2019) and empirically verify that the amount of inverted elements introduced in the piece-wise map gracefully decays to zero for growing numbers of eigenfunctions (Figure 7). On average, already using $5\%$ of the spectrum the number of flipped tetrahedra is below $5\%$. It goes below $1\%$ if $20\%$ of the spectrum is considered (Table 1). We consider both the LBO spectrum and its extension equipped with three additional functions that encode the extrinsic embedding (CMH). As shown in (Marin et al., 2019; Melzi et al., 2020) CMH is better suited to transfer discrete signals such as the surface mesh connectivity. This advantage also manifests on volumes, as reflected in the second and fourth columns in Figure 7 and in the second and fourth lines of Table 1. For efficiency, the average running times for computing the larger bases are reported in the last column of the same table. As can be noticed the extrapolation algorithm is roughly 2×2$\times$2 × faster than volumetric transfer, because in this case only the volumetric spectrum of a single mesh must be computed. Interestingly, if the mesh resolution is sufficiently low and the computation of a higher percentage of the spectrum is computationally feasible, we can directly devise a fully injective piecewise linear map, as shown in Figure 9, where $25\%$ of the spectrum is considered.
When computational resources are limited and the full spectrum cannot be computed, volumetric functional mapping still provides a valuable initial solution for untangling algorithms such as (Garanzha et al., 2021; Du et al., 2020), which are then required to remove the few inverted tetrahedra present in the map. This is a standard procedure for the generation of injective linear maps, where mesh untangling is typically bootstrapped using the Tutte embedding, which is known to flip volumetric cells (Alexa, 2023). Volumetric functional map provides a better initial solution than Tutte. To prove this, we considered the 1405 failures produced in the large-scale validation described in (Nigolian et al., 2023, §6.3). Substituting the initial solution provided by Tutte with the one produced with our framework, we were able to produce fully bijective maps in the 85% of their failure cases using both LBO and CMH basis functions, thus dramatically increasing the success rate of this pipeline. Two representative examples are shown in Figure 8.
5.3. Segmentation Transfer
The functional maps framework proved to be particularly successful in the area of statistical shape analysis (Stegmann and Gomez, 2002). In particular, the study of medical data primarily benefits from accurate solutions for analyzing large collections of shapes (Heimann and Meinzer, 2009), and the literature proves that the functional maps framework can be successfully applied for medical applications (Melzi et al., 2016; Magnet et al., 2023; Maccarone et al., 2024). A standard approach for analyzing collections of shapes is to determine or build a template surface which acts as a median shape for comparing the entire dataset (Huang et al., 2019). However, existing approach are limited to the analysis of surfaces, while often medical data are better represented as 3D volumes, such in the case of MRI scans (Rinck, 2019).
@@ 5.4. Shape Matching
We consider both the outer surfaces of our reference ground truth volumetric dataset (which we will refer to as VOL) (Su et al., 2019) and well established datasets for shape matching, such as SHREC’19 (Melzi et al., 2019a), SHREC’20 (Dyke et al., 2020), and TOPKIDS (Lähner et al., 2016), which we all tetrahedralize using fTetWild (Hu et al., 2020). Overall, these datasets stress shape matching algorithms in different ways, exhibiting geometric and topological noise (SHREC’19), strongly non-isometric matches (SHREC’20) and topological changes (SHREC’20 and TOPKIDS). For SHREC’20, we discard two partial meshes because the inside/outside system of fTetWild was not able to correctly reproduce a volumetric mesh that allows a meaningful comparison with surface methods. More discussion about the tetrahedralization of the surfaces is reported in Appendix E.
@@ 6. Conclusions
First and foremost, a theoretical study of what properties are invariant under a volume functional map must be carried out. Despite our empirical studies show that volumetric maps are more accurate than their surface counterparts, a more profound theoretical analysis of this behavior would provide useful insights for the development of new solutions.
@@
Appendix A Choice of the Laplacian
The original functional maps paper introduced the use of the LBO eigenbasis because it satisfies the desired properties of compactness of representation and stability under near-isometric deformations (Ovsjanikov et al., 2012, §5.1). Furthermore, later works from Aflalo et al. (2015; 2016) have shown that, in a general setting, and especially when working with smooth signals, the LBO eigenbasis of a manifold $\mathcal{M}$ (of arbitrary dimension) is the optimal orthonormal basis for representing a finite-dimensional subspace of $L^{2}(\mathcal{M})$. In this setting, it is reasonable to extend the choice of the LBO to the volumetric domain, especially considering that volumetric deformations do not induce a larger geodesic distortion with respect to near-isometric surface deformations (see Appendix C for an empirical study in this regard).
On the other hand, it should be noted that discretizing the LBO operator in a way that fulfills all the algebraic properties of the same operator in the smooth setting is often difficult (Wardetzky et al., 2007). In literature there exist multiple alternative discretizations (Jacobson, 2024; Alexa et al., 2020; Mullen et al., 2011), each one with its own pros and cons. Among the various alternatives, we chose the cotangent formula due to its popularity, ease of implementation, and efficiency. This choice is not restrictive. Any alternative discretization can be used equivalently. Nonetheless, since most applications that we address involve analyzing information at the surface (which is the boundary of a tetrahedral mesh), we are forced to a discretization that imposes Neumann’s boundary conditions to avoid zero-valued eigenfunctions on the surface.
Volumetric Cotangent Formula
We discretize the LBO of a tetrahedral mesh $\mathcal{M}=(V_{\mathcal{M}},T_{\mathcal{M}})$ with the following per vertex relation
(11) $\Delta(v_{i})=\sum_{j\in N(v_{i})}\!\!w_{ij}(v_{i}-v_{j})\,,$
using the $n$-dimensional cotangent formula (Liao et al., 2009; Crane, 2019) to compute per edge weights $w_{ij}$. These weights consider all tetrahedra $ijkl$ incident to edge $ij$, according to the following formula
(12) $w_{ij}=\frac{1}{6}\sum_{ijkl}|v_{k}-v_{l}|\cot\theta_{kl}\,,$
where angle $\theta_{kl}$ denotes the dihedral angle at the edge $kl$ opposite to edge $ij$ w.r.t. tetrahedron $ijkl$.
Appendix B Spectral Comparison
The eigenvalues of the Laplace-Beltrami operator form a non-decreasing sequence that grows with a rate $\lambda_{k}\in\mathcal{O}(\sqrt[d]{k^{2}})$, where $d$ is the dimensionality of the manifold (Weyl, 1911). A direct comparison of the eigenvalues is thus misleading, as the difference will be dominated by the eigenvalues with larger index. One possibility for handling this issue is to introduce relative differences in the comparison. Alternatively, Moschella et al. (2022) introduced the offset representation for comparing and handling Laplacian spectra. Given two manifolds $\mathcal{M}$ and $\mathcal{N}$, the offset difference can be defined as follows
(13) $\mathrm{offset}(\bm{\lambda}_{\mathcal{M}},\bm{\lambda}_{\mathcal{N}})=|\mathrm{off}(\bm{\lambda}_{\mathcal{M}})-\mathrm{off}(\bm{\lambda}_{\mathcal{N}})|\,,$
(14) $\mathrm{off}(\bm{\lambda}_{\mathcal{M}})=(\mathrm{off}(\lambda_{\mathcal{M},1}),\,\mathrm{off}(\lambda_{\mathcal{M},2}),\,\cdots\,)\,,$
(15) $\mathrm{off}(\lambda_{\mathcal{M},i})=\lambda_{\mathcal{M},i}-\lambda_{\mathcal{M},i-1}\,.$
This method for comparing eigenvalues is designed to compensate for the Weyl’s law on 2-dimensional manifolds: since $\lambda_{k}\in\mathcal{O}(k)$, comparing consecutive differences (which asymptotically behave like a constant function) is more informative than comparing absolute values. However, when dealing with 3-dimensional manifolds, the Weyl’s law changes to $\lambda_{k}\in\mathcal{O}(\sqrt[3]{k^{2}})$ and such consecutive differences are not asymptotically constant anymore. This change to the asymptotic behavior also affects the relative difference: if we consider that the $k$-th eigenvalue can be written as the Weyl’s asymptotic relation plus a correction term, namely $\lambda_{k}=c\sqrt[d]{k^{2}}+r$, we have that
[…]
As a consequence neither the relative difference or the offset difference have the same scale for 2-dimensional and 3-dimensional manifolds.
However, we can use the Weyl’s law to compensate the non-linear growth of the eigenvalues. Since $\lambda_{k}\in\mathcal{O}\sqrt[3]{k^{2}}$, before computing the offsets we can transform the eigenvalues as $\lambda_{k}^{\prime}=\sqrt{\lambda_{k}^{3}}\in\mathcal{O}(k)$. With this transformation, the error measures has the same scale for both the surfaces and the volumes. Clearly, this holds under the assumption that all the 2-dimensional manifolds have unitary area and all the 3-dimensional manifolds have unitary volume, which is easy to enforce as a preprocessing step.
Appendix C Volumetric Isometries
Figure 13. For each subject, the relative geodesic distortion between pairs of volumes (orange) and pairs of surfaces (blue). The plot provides both the average distortion and the standard deviation.
The functional maps framework relies on the assumption that the two shapes are isometric or quasi-isometric. In practice, the framework proposed by Ovsjanikov et al. (2012) works well even under mild non-isometric deformations and meshes at a different scale. However, when dealing with strong non-isometric distortion, more sophisticated approaches are needed, which cannot be easily lifted to manifolds of higher dimensions (Panine et al., 2022).
In order to verify that the functional maps framework can be meaningfully transported to a volumetric setting, we need to verify that the level of isometric distortion does not change when a 3D shape is represented as a 2-dimensional surface or a 3-dimensional volume. The dataset introduced by Su et al. (2019) contains 40 pairs of tetrahedral meshes in bijective correspondence, and thus it is easy to compute the distortion of geodesic distances induced by the correspondence. For each pair of shapes $\mathcal{M}$ and $\mathcal{N}$, we consider 100k random pairs of points $x,y\in\partial\mathcal{M}$. We then compute the geodesic distortion induced by the ground truth correspondence $\pi:\mathcal{M}\to\mathcal{N}$ for both the surface and the volumetric setting, respectively as
(17) $\mathrm{SurfErr}(x,y)=\frac{|d_{\partial\mathcal{M}}(x,y)-d_{\partial\mathcal{N}}(\pi(x),\pi(y)|}{d_{\partial\mathcal{M}}(x,y)}\,,$
(18) $\mathrm{VolErr}(x,y)=\frac{|d_{\mathcal{M}}(x,y)-d_{\mathcal{N}}(\pi(x),\pi(y)|}{d_{\mathcal{M}}(x,y)}\,,$
where $d_{\mathcal{S}}$ denotes the approximation of the geodesic distance in the volume $\mathcal{S}$ provided by the Dijkstra distance, and $d_{\partial\mathcal{S}}$ denotes the analogous approximation over the boundary surface $\partial\mathcal{S}$.
The results are summarized in Figure 13, where we report for each pair of shapes the average distortion over the surface (in blue) and the volume (in orange), together with the standard deviation. We can see that, in all cases, the distortions are comparable, with the tetrahedral meshes generally achieving a slightly better preservation of the geodesic distances. As a consequence, we guess that, for most practical applications, the property of a pair of shapes of being isometric is preserved between a surface and volumetric representation, thus allowing us to safely extend the functional maps framework from surfaces to volumes.
Appendix D Boundary Trace of Eigenfunctions
Figure 14. For a varying basis size, we try to reconstruct the LBO eigenfunctions of the surface using the restriction of the volumetric LBO eigenfunctions at the boundary. The reconstruction error lies in the range $[0,1]$.
As mentioned in Section 4.2, given the volumetric LBO eigenfunctions $\Phi_{\mathcal{M}}$ of a 3-dimensional manifold $\mathcal{M}$, the restriction $\Phi_{\partial\mathcal{M}}$ of such basis at the boundary surface $\partial\mathcal{M}$ (often also called boundary trace) forms a set of linearly independent functions over $\partial\mathcal{M}$ (Lasiecka and Triggiani, 1983; Triggiani, 2008).
However, $\Phi_{\partial\mathcal{M}}$ is not necessarily equivalent to the LBO eigenbasis $\mathbf{\Psi}_{\partial\mathcal{M}}$ of the surface, meaning that it could be a sub-optimal basis (Aflalo et al., 2015, 2016). In order to evaluate the expressive power of the boundary trace of the volumetric eigenfunctions, we try to reconstruct the basis $\mathbf{\Psi}_{\partial\mathcal{M}}$ using the basis $\Phi_{\partial\mathcal{M}}$. For our experiment, we use the 80 shapes from the dataset presented by Su et al. (2019) and we try to reconstruct the first $k$ eigenfunctions of the surface with the first $k$ boundary traces. In particular, we first compute the reconstruction $\mathbf{\Psi}^{\prime}=\Phi_{\partial\mathcal{M}}\Phi_{\partial\mathcal{M}}^{\dagger}\mathbf{\Psi}_{\partial\mathcal{M}}$. Then, for each surface eigenfunction $\psi_{i}$ and its reconstruction $\psi_{i}^{\prime}$, we compute the reconstruction error as
(19) $\mathrm{err}(\psi_{i})=\int_{\partial\mathcal{M}}(\psi_{i}(x)-\psi_{i}^{\prime}(x))^{2}\,\mathrm{d}x\,.$
This error measure lies in the range $[0,1]$.
In Figure 14 we show the distributions of the reconstruction errors across the entire dataset for different values of $k$. As shown in the figure, very few of the surface eigenfunctions are close to be orthogonal to the boundary traces. However, the boundary restriction of the volumetric eigenfunctions is able to faithfully represent a significant portion of the surface eigenbasis, and provides a reasonable approximation for most of them. Therefore, while not being the optimal basis, the boundary trace of the volumetric LBO eigenfunctions still provides a good candidate for representing a $k$-dimensional function space over the surface.
Appendix E Volumetric Shape Matching Datasets
As mentioned in Section 5.4, well-established datasets for shape matching only provide surfaces. On top of that, most of the available recent datasets try to highlight some challenges in the problem of finding a correspondence; this includes partialities, holes, heavy topological noise, self-intersections, clutter, and so on.
Dropping the assumptions of manifoldness, watertightness, and absence of self-intersections for the surfaces, the most robust tool for producing tetrahedral meshes results to be fTetWild (Hu et al., 2020). Nonetheless, for some of the surfaces, the problem of finding a tetrahedralization is so ill-posed that any resulting volume would not provide meaningful information about the 3D shape.
Indeed, for SHREC’20, we discard the two partial meshes shown in Figure 15 because the inside/outside system of fTetWild was not able to correctly reproduce a volumetric mesh that allows a meaningful comparison with surface methods. For the same reason, we are not able to consider alternative datasets such as SMAL (Zuffi et al., 2017), which contains numerous self-intersections that yield topological changes during tetrahedralization. Attempts to sanitize these models with various combinations of geometry processing tools (Portaneri et al., 2022; Hang, 2015; Livesu, 2019; Diazzi and Attene, 2021; Diazzi et al., 2023) ended with no luck. Indeed, our difficulties in creating a dataset that allows to meaningfully compare surface-based and volume-based shape matching methods raises an open research challenge that will hopefully be addressed by the community in the near future (Section 6).